40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 21 09:53:46 2013, Annalisa, Update, Locking, Finding the beat note 
    Reply  Thu Mar 21 10:19:28 2013, Koji, Update, Locking, Finding the beat note 
       Reply  Thu Mar 21 13:11:42 2013, Annalisa, Update, Locking, Finding the beat note Beat_note_setup.jpg
          Reply  Thu Mar 21 13:50:08 2013, Jamie, Update, Locking, incident angles 
             Reply  Thu Mar 21 14:02:45 2013, Annalisa, Update, Locking, incident angles 
                Reply  Fri Mar 22 23:23:38 2013, Annalisa, Update, Locking, Beat note still missing Beat_note_-_new_setup.jpg
Message ID: 8327     Entry time: Thu Mar 21 13:11:42 2013     In reply to: 8323     Reply to this: 8328
Author: Annalisa 
Type: Update 
Category: Locking 
Subject: Finding the beat note 

Quote:

Give us more info on the elog:
What PD are you using? How much power the beams on the recombining BS are? What kind of BS is it?
How are you looking for the beat note? (on the scope? or spectrum analyzer?)
What was the scanned temp range?

Three points to be checked:

- Polarization

- Alignment

- Temperature

Experimental Setup

I'm using a 1611 New Focus PD (1 GHZ, with maximum input power 1mW), and the total power hitting on the PD is of about 0.650 mW.

The current of the NPRO laser is set to 1.38 A, so that the input power is 19 mW. The beam is initially damped by a 10% reflection BS and then it hits a 33% reflection BS (where it recombines with the PSL pick-off beam) with 2 mW power.

After this second BS the power is reduced to 0.592 mW.

The PSL pick-off hits on the 33% reflection BS with 65.5 uW power, and it exit with a 47 uW power.

I connected a power supply to apply a Voltage to the slow frequency BNC, in way to tune the laser frequency.

I'm using the AGILENT 4395A Spectrum analyzer to make the measurement. I tried to use the HEWELETT PACKARD 8591E spectrum analyzer, but the monitor didn't turn on.

The temperature spanned until now in only of about 10 deg C, because I realized that I needed a better alignment, so I added a lens in front of the PD and I did a better alignment. 

Moreover, the current of the laser is too low, so I need to increase it and add more beam splitters in the beam path to dump the beam, in way to don't reach the PD threshold.

I knew that both the beams are s-polarized, but maybe I can check it again.

Attachment 1: Beat_note_setup.jpg  151 kB  | Hide | Hide all
Beat_note_setup.jpg
ELOG V3.1.3-